How do you find the absolute value of an open sentence when it has parenthesis? Ex: |6(3d-5)| is less than or equal to 14
Can you help me?

Answer by stanbon(29602)

You have to solve two equations and then combine their answers, as follows:
6(3d-5)<=14
3d-5<= 7/3
3d<=7/3 + 15/3
d<= 22/9
-6(3d-5)<=14
-18d+30<=14
-18d<=-16
d>=8/9
Combine the two answers to get:
8/9<=d<=22/9
Cheers,
Stan H.